1 + 2/3 divided by 2 + 1/2
Answer by checkley75(3666)

1+2/3=(3+2)/3=5/3
2+1/2=(4+1)/2=5/2
NOW TO DIVIDE FRACTIONS THE DENOMINATOR FRACTION NEEDS TO BE INVERTED & THE TWO FRACTIONS ARE THEN MULTIPLIED THUS:
5/3*2/5=10/15=2/3 ANSWER.